OpenStudy (anonymous):

Find lim_{x rightarrow infty} (sqrt{x^2+5x+1}-x) [hint: rationalize.]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(\sqrt{x^2+5x+1}-x)\]

OpenStudy (michele_laino):

hint: we can write: \[\large \sqrt {{x^2} + 5x + 1} - x = \frac{{{x^2} + 5x + 1 - {x^2}}}{{\sqrt {{x^2} + 5x + 1} + x}}\]

OpenStudy (michele_laino):

more explanation: \[\large \begin{gathered} \sqrt {{x^2} + 5x + 1} - x = \frac{{\left( {\sqrt {{x^2} + 5x + 1} - x} \right)\left( {\sqrt {{x^2} + 5x + 1} + x} \right)}}{{\sqrt {{x^2} + 5x + 1} + x}} = \hfill \\ = \frac{{{x^2} + 5x + 1 - {x^2}}}{{\sqrt {{x^2} + 5x + 1} + x}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Yes, I got that part

OpenStudy (michele_laino):

ok! Then, we have to factor out x at numerator and at denominator, so we get: \[\frac{{5x + 1}}{{\sqrt {{x^2} + 5x + 1} + x}} = \frac{{x\left( {5 + 1/x} \right)}}{{x\left( {\sqrt {1 + 5/x + 1/{x^2}} + 1} \right)}} = \frac{{5 + 1/x}}{{\sqrt {1 + 5/x + 1/{x^2}} + 1}}\]

OpenStudy (michele_laino):

that expression is correct if \[\large x \to + \infty \]

OpenStudy (anonymous):

oh okay

OpenStudy (michele_laino):

whereas if \[x \to - \infty \] we have: \[\frac{{5x + 1}}{{\sqrt {{x^2} + 5x + 1} + x}} = \frac{{x\left( {5 + 1/x} \right)}}{{ - x\left( {\sqrt {1 + 5/x + 1/{x^2}} + 1} \right)}} = - \frac{{5 + 1/x}}{{\sqrt {1 + 5/x + 1/{x^2}} + 1}}\]

OpenStudy (anonymous):

Just wondering where did you get the 1/x^2 in the denominator?

OpenStudy (michele_laino):

since we have the subsequent steps: \[\large \sqrt {{x^2} + 5x + 1} + x = \sqrt {{x^2}\left( {1 + \frac{5}{x} + \frac{1}{{{x^2}}}} \right)} + x = \left| x \right|\sqrt {1 + \frac{5}{x} + \frac{1}{{{x^2}}}} + x\]

OpenStudy (michele_laino):

\[\large \begin{gathered} \sqrt {{x^2} + 5x + 1} + x = \sqrt {{x^2}\left( {1 + \frac{5}{x} + \frac{1}{{{x^2}}}} \right)} + x = \hfill \\ = \left| x \right|\sqrt {1 + \frac{5}{x} + \frac{1}{{{x^2}}}} + x \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

please, I have made an error, when \[x \to - \infty \] the original expression, can be rewritten as follows: \[\frac{{5x + 1}}{{\sqrt {{x^2} + 5x + 1} + x}} = \frac{{x\left( {5 + 1/x} \right)}}{{x\left( { - \sqrt {1 + 5/x + 1/{x^2}} + 1} \right)}} = \frac{{5 + 1/x}}{{ - \sqrt {1 + 5/x + 1/{x^2}} + 1}}\]

OpenStudy (anonymous):

Oh I see... Then when an sub in the infinity and follow limit law and get rid of the 1/x, 5/x and 1/x^2 beacuse they will equal to zero, right? So we get 5/sqrt 2? @Michele_Laino

OpenStudy (anonymous):

I mean 5/2

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

YAY! Thank you so much! :)

OpenStudy (michele_laino):

Thank you! :)